Solve for h if 3f + 2.5h= 240. The answer in my math book says h= -3f+96/2.5 but my answer is h= -3f+96. Which is correct?
Thanks Alessia

Answer by checkley79(3341)

3f + 2.5h= 240
2.5h=240-3f
h=(240-3f)/2.5 divide both terms in the numerator by 2.5
h=96-1.2f ans.